# Divide and Conquer: High-Resolution Industrial Anomaly Detection via Memory Efficient Tiled Ensemble

### 07 Mar 2024 v1

@@ Abstract
Industrial anomaly detection is an important task within computer vision with a wide range of practical use cases. The small size of anomalous regions in many real-world datasets necessitates processing the images at a high resolution. This frequently poses significant challenges concerning memory consumption during the model training and inference stages, leaving some existing methods impractical for widespread adoption. To overcome this challenge, we present the tiled ensemble approach, which reduces memory consumption by dividing the input images into a grid of tiles and training a dedicated model for each tile location. The tiled ensemble is compatible with any existing anomaly detection model without the need for any modification of the underlying architecture. [^1] Available on GitHub:
https://github.com/openvinotoolkit/anomalib
By introducing overlapping tiles, we utilize the benefits of traditional stacking ensembles, leading to further improvements in anomaly detection capabilities beyond high resolution alone. We perform a comprehensive analysis using diverse underlying architectures, including Padim, PatchCore, FastFlow, and Reverse Distillation, on two standard anomaly detection datasets: MVTec and VisA. Our method demonstrates a notable improvement across setups while remaining within GPU memory constraints, consuming only as much GPU memory as a single model needs to process a single tile. [^2] Research conducted during Google Summer of Code 2023 at OpenVINO (Intel)
@@ 1 Introduction
Tiling mechanisms, which subdivide the input images into a rectangular grid of tiles as a pre-processing step, have been used to process images at a high resolution while keeping memory use low [25]. By passing individual tiles to the model as input instead of full images, tiling reduces the model’s input dimensions, while maintaining the effective input resolution of the images content-wise. However, this approach may not always be ideal, particularly for methods sensitive to object alignment [23], as using a single model for all patches may compromise spatial information preservation.
@@ 3 Method
The tiled ensemble method is structured as a series of sequential steps. The method initially divides the high-resolution image into tiles (Section 3.1), followed by training individual models for each tile location (Section 3.2). Once predictions are obtained, a merging mechanism is utilized to produce full-image-level data, followed by standard post-processing steps (Section 3.3).
@@ 3.2 Training and inference
Formally, a separate model $M_{i,j}$ is trained for each tile location for the entire set of tiles $\mathcal{T}$, defined in Equation 1, resulting in a set of models:
@@ 3.3 Merging
The borders of the tiles create seams, leading to undesirable disturbances in the image. To mitigate this issue and enhance the outcomes, a Gaussian filter is applied for smoothing. This smoothing process is confined to a narrow region surrounding the seam, as depicted in Figure 3. The default width of the smoothing region is $10\%$ of the tile width on each side of the seam. Since the output at this stage is the same as in a standard anomaly detection model, it is fully compatible with any of the common post-processing steps, such as normalization and thresholding.
@@ 4.1 Experimental details
Datasets. The method is evaluated on two established industrial datasets: MVTec AD [6] and VisA [46]. MVTec AD and VisA datasets comprise 15 and 12 categories, respectively. Each category consists of a training set containing only normal images and a test set, comprising both normal and anomalous images, with their corresponding pixel-precise ground truth annotations. The anomalies vary in types, shapes, and scales, with the prevalence of larger anomalies in MVTec AD and smaller defects in VisA.
@@ 4.2 Evaluation setups
Common properties. Each setup is trained on every category, with every run repeated 3 times using a different random seed. Center-crop is disabled for all methods, and a consistent batch size of 32 is used, except for Patchcore where it’s set to 8 due to memory limitations. The backbone used in all setups is ResNet18, to keep the comparison fair. Following [4, 19] FastFlow, and Reverse Distillation are limited to 200 steps for all setups. Other properties are kept the same as provided by the original authors of the models and as implemented in Anomalib [2].
@@ 5 Results
Results on VisA with all 6 setups are displayed in Table 2. A tiled ensemble with overlapping tiles (ENS9) achieves the best results in terms of anomaly detection for Padim, FastFlow, and Reverse Distillation, significantly outperforming baseline single model (SM256) and single model processing the same resolution of $512\times 512$ (SM512) in all three cases. It also achieves the best results for anomaly localization for FastFlow and Reverse Distillation.
@@
Figure 4: Effects of increasing resolution for Padim on VisA PCB3 category. The first column contains full and zoomed-in input images with corresponding ground truth. The next columns depict full and zoomed-in anomaly maps with their corresponding binary segmentation. The resolution of results is written above each block. Notice how localization improves with higher resolution. The tiled ensemble uses a setup with $h^{t}=w^{t}=256$, $s_{h}=s_{w}=256$
@@
More detailed results for each category and setup with included standard deviation on MVTec AD and VisA are included in Appendix B.
@@
Figure 6: Localization results in terms of AUPRO on VisA PCB3 category for a single model and tiled ensemble with different resolutions. The corresponding memory consumption of each setup is shown on the right. The tiled ensemble uses a setup with $h^{t}=w^{t}=256$, $s_{h}=s_{w}=256$
@@
Figure 7: Effect of dataset size on anomaly detection and localization performance of the tiled ensemble (ENS9) and single model with $256\times 256$ resolution (SM256). Each point represents a single dataset category from MVTec AD or VisA. Trend lines and confidence intervals added for interpretability. X-axis: average number of anomalous pixels per defective image relative to image size. Y-axis: average performance of setup across all four model architectures.
Latency and throughput. The latency and throughput are presented in Figure 8.
@@
Primarily, the latency overhead stems from the time needed to transfer individual models to GPU and back, which does not affect throughput as significantly since the model’s time on GPU is better utilized. For some models like Patchcore, the throughput of a 4-tiled ensemble exceeds that of a single model processing an equivalent resolution. The preliminary studies showed that the time needed for tiled ensemble inference on GPU still outperforms the inference of a single model with increased resolution on CPU in terms of latency by around 4 times, and in terms of throughput by around 80 times.
@@ 6 Conclusion
This paper introduces a tiled ensemble approach to effectively detect and localize small anomalies in high-resolution images. This has been a challenge due to the high GPU memory demands required by existing approaches. The method works by dividing the image into smaller tiles and training a dedicated model for each tile location. This strategy ensures that the GPU memory usage remains comparable to that of a single model that processes an image the size of one tile. By employing overlapping tiles, this approach also takes advantage of the performance improvements associated with traditional stacking ensemble methods, which further improve performance compared to those achievable by simply increasing image resolution.
@@
Limitations. Despite its promising statistical results, our approach has a notable inference overhead, which can be partially mitigated through batched inference. This can be a reasonable sacrifice in cases where the resolution is very large, to enable detection with resolutions that previously were not feasible. Heckler et al. [19] suggest and Heckler. and König. [18] verify that in certain scenarios, strategically choosing a single layer can outperform an ensemble of multiple layers and backbones. This insight could complement our approach. It potentially opens avenues for future research where each separate tile location uses the most suitable layer, further optimizing anomaly detection in high-resolution images. The approach has additional limitations due to the tiling of the input, potentially preventing accurate logical anomaly detection as tiling prevents efficient capturing of the global context.
@@ Appendix A Anomaly scales per category
Figure 9 shows the average anomalous pixel ratio (what percentage of defective image is covered by the defect) present in test set of each category from MVTecAD [6] and VisA [46]. The ratio is calculated on all defective images from the category with a resolution of $512\times 512$. VisA categories contain notably smaller defect, especially in categories such as candle, macaroni 1, and macaroni 2, where the anomalous pixels on average cover less than $1\%$ of image.
@@
Appendix B Results of all categories for each setup
This section contains tables with category specific results for all architectures and their setups.
@@ Appendix B Results of all categories for each setup
Appendix C Additional qualitative examples.

### 08 Apr 2024 v2

@@ Abstract
Industrial anomaly detection is an important task within computer vision with a wide range of practical use cases. The small size of anomalous regions in many real-world datasets necessitates processing the images at a high resolution. This frequently poses significant challenges concerning memory consumption during the model training and inference stages, leaving some existing methods impractical for widespread adoption. To overcome this challenge, we present the tiled ensemble approach, which reduces memory consumption by dividing the input images into a grid of tiles and training a dedicated model for each tile location. The tiled ensemble is compatible with any existing anomaly detection model without the need for any modification of the underlying architecture. By introducing overlapping tiles, we utilize the benefits of traditional stacking ensembles, leading to further improvements in anomaly detection capabilities beyond high resolution alone. We perform a comprehensive analysis using diverse underlying architectures, including Padim, PatchCore, FastFlow, and Reverse Distillation, on two standard anomaly detection datasets: MVTec and VisA. Our method demonstrates a notable improvement across setups while remaining within GPU memory constraints, consuming only as much GPU memory as a single model needs to process a single tile.  [^1] Available as part of Anomalib:
https://github.com/openvinotoolkit/anomalib.
 [^2] Research conducted during GSoC 2023 at OpenVINO.
@@ 1 Introduction
Tiling mechanisms, which subdivide the input images into a rectangular grid of tiles as a pre-processing step, have been used to process images at a high resolution while keeping memory use low [25]. By passing individual tiles to the model as input instead of full images, tiling reduces the model’s input dimensions, while maintaining the effective input resolution of the images content-wise. This approach may not always be ideal, particularly for methods sensitive to object alignment [23], as using a single model for all patches may compromise spatial information preservation.
@@ 3 Method
The tiled ensemble method is structured as a series of sequential steps. The method initially divides the high-resolution image into tiles (Section 3.1), followed by training individual models for each tile location (Section 3.2). Once predictions are obtained, a merging mechanism is utilized to produce full-image-level data (Section 3.3), followed by standard post-processing steps.
@@ 3.2 Training and inference
Formally, a separate model $M_{i,j}$ is trained for each tile location in the entire set of tiles $\mathcal{T}$, defined in Equation 1, resulting in a set of models:
@@ 3.3 Merging
The borders of the tiles create seams, leading to undesirable disturbances in the image. To mitigate this issue and enhance the outcomes, a Gaussian filter is applied for smoothing. This smoothing process is confined to a narrow region surrounding the seam, as depicted in Figure 3. The default width of the smoothing region is $10\%$ of the tile width on each side of the seam. In line with standard anomaly detection procedures, the final classification and localization predictions can be obtained by applying a thresholding mechanism to the image-level anomaly scores and anomaly maps respectively.
@@ 4.1 Experimental details
Datasets. The method is evaluated on two established industrial datasets: MVTec AD [6] and VisA [46]. MVTec AD and VisA datasets comprise 15 and 12 categories, respectively. Each category consists of a training set containing only normal images and a test set, comprising both normal and anomalous images, with their corresponding pixel-precise ground truth annotations. The anomalies vary in types, shapes, and scales, with the prevalence of larger anomalies in MVTec AD and smaller defects in VisA. An analysis of defect scales is presented in Appendix A.
@@ 4.2 Evaluation setups
Common properties. Each setup is trained on every category, with every run repeated 3 times using a different random seed. A consistent batch size of 32 is used, except for Patchcore where a batch size of 8 is used due to memory limitations. The backbone used in all setups is ResNet18, to keep the comparison fair. Following [4, 19] FastFlow, and Reverse Distillation are limited to 200 steps for all setups. Other properties are kept the same as provided by the original authors of the models and as implemented in Anomalib [2].
@@ 5 Results
Results on VisA with all 6 setups are displayed in Table 2. A tiled ensemble with overlapping tiles (ENS9) achieves the best anomaly detection results for Padim, FastFlow, and Reverse Distillation, significantly outperforming baseline single model (SM256) and single model processing the same resolution of $512\times 512$ (SM512) in all three cases. ENS9 also achieves the best anomaly localization results for FastFlow and Reverse Distillation.
@@
Figure 4: Effects of increasing resolution for Padim on VisA PCB3 category. The first column contains full and zoomed-in input images with corresponding ground truth. The next columns depict full and zoomed-in anomaly maps with their corresponding binary segmentation. The resolution of results is written above each block. Notice how localization improves with higher resolution. The tiled ensemble uses a setup with $h^{t}=w^{t}=256$, $s_{h}=s_{w}=256$.
@@
More detailed results for each category and setup with included standard deviation on MVTec AD and VisA are included in Appendix D.
@@
Figure 6: Localization results in terms of AUPRO on VisA PCB3 category for a single model and tiled ensemble with different resolutions. The corresponding memory consumption of each setup is shown on the right. The tiled ensemble uses a setup with $h^{t}=w^{t}=256$, $s_{h}=s_{w}=256$.
@@
Figure 7: Effect of defect size on anomaly detection and localization performance of the tiled ensemble (ENS9) and single model with $256\times 256$ resolution (SM256). Each point represents a single dataset category from MVTec AD or VisA. Trend lines and confidence intervals added for interpretability. X-axis: average number of anomalous pixels per defective image relative to image size. Y-axis: average performance of setup across all four model architectures.
Latency and throughput. The latency and throughput are presented in Figure 8. The low throughput and high latency of ENS4 and ENS9 can be attributed to the increased computational complexity of these setups (Appendix B) and show that the performance advantage of the tiled ensemble comes at the cost of an increased runtime. The latency overhead likely stems from the time needed to transfer individual models to GPU and back, which does not affect throughput as significantly since the model’s time on GPU is better utilized. For some models like Patchcore, the throughput of a 4-tiled ensemble exceeds that of a single model processing an equivalent resolution. The preliminary studies showed that the time needed for tiled ensemble inference on GPU still outperforms the inference of a single model with increased resolution on CPU in terms of latency by around 4 times, and in terms of throughput by around 80 times. The training time of all setups is presented in Appendix C.
@@ 6 Conclusion
This paper introduces a tiled ensemble approach to effectively detect and localize small anomalies in high-resolution images, which has been a challenge due to the high GPU memory demands required by existing approaches. The tiled ensemble approach addresses this by dividing the image into smaller tiles and training a dedicated model for each tile location. This strategy ensures that the GPU memory usage remains comparable to that of a single model that processes an image the size of one tile. By employing overlapping tiles, the tiled ensemble also takes advantage of the performance improvements associated with traditional stacking ensemble methods, which further improve performance compared to those achievable by simply increasing image resolution.
@@
Limitations. Despite its promising statistical results, our approach has a notable latency overhead, which can be partially mitigated through batched inference. This can be a reasonable sacrifice in cases where the resolution is very large, to enable detection with resolutions that previously were not feasible. As suggested by Heckler et al. [19] and verified by Heckler. and König. [18], strategically choosing a single layer can outperform an ensemble of multiple layers and backbones in certain scenarios. Building on this insight, future research should investigate whether selecting the most suitable layer, or set of layers, for each tile location could further optimize anomaly detection in high-resolution images. Finally, the experiments of the current study did not cover logical anomaly detection benchmarks, which could potentially suffer from a loss of global context as a result of processing each tile location separately.
@@ Appendix A Anomaly scales per category
Figure 9 shows the average anomalous pixel ratio (what percentage of defective image is covered by the defect) present in the test set of each category from MVTecAD [6] and VisA [46]. The ratio is calculated on all defective images from the category with a resolution of $512\times 512$. VisA categories contain notably smaller defects, especially in categories such as candle, macaroni 1, and macaroni 2, where the anomalous pixels on average cover less than $1\%$ of the image.
@@
Appendix B Parameter sizes
Table 3 contains the parameter count (in millions) for each setup and architecture. Since the tiled ensemble consists of unchanged underlying architectures, the parameter sizes are increased by the factor equivalent to the number of models in an ensemble.
Setup PatchCore Padim FastFlow Reverse Distillation
SM256 2.8 2.8 9.7 18.2
SM512 2.8 2.8 12.5 18.2
ST4 2.8 2.8 9.7 18.2
ST9 2.8 2.8 9.7 18.2
ENS4 11.1 11.1 38.9 72.7
ENS9 25.0 25.0 87.6 163.5
Table 3: Parameters (million) for each architecture and each setup.
Appendix C Training time
Table 4 contains the training duration of each architecture with each setup. The results are averaged across all categories from MVTecAD [6] and VisA [46] and 3 runs with different seeds. In cases of FastFlow and Reverse Distillation, the training time is approximately extended by a factor equivalent to the number of models in an ensemble. This is expected due to the cumulative increase in epochs, while the training workload of each model inside the ensemble remains in line with a single model processing a $256\times 256$ resolution.
In the case of PatchCore and Padim, this does not hold since they are not trained using backpropagation. Therefore, the time required for training doesn’t scale equivalently when directly increasing resolution (from SM256 to SM512) or when achieving higher processed resolution through multiple smaller models operating within an ensemble (from SM256 to ENS4/ENS9).
Setup PatchCore Padim FastFlow Reverse Distillation
SM256 213.1 772.0 556.3 721.5
SM512 2273.0 2423.5 799.6 838.8
ST4 2254.6 2400.7 805.1 836.5
ST9 2275.2 2423.0 851.9 867.5
ENS4 879.9 444.9 2561.8 2970.9
ENS9 1810.6 974.6 5857.7 6876.8
Table 4: Training time in seconds for each architecture and each setup. Results are averaged over all categories and 3 runs with different seeds.
Appendix D Results of all categories for each setup
This section contains tables with category-specific results for all architectures and their setups.
@@ Appendix D Results of all categories for each setup
Appendix E Additional qualitative examples.